# Weighted Missing Linear Discriminant Analysis: An Explainable Approach for Classification with Missing Data

### v1

@@ Abstract
As Artificial Intelligence (AI) models are gradually being adopted in real-life applications, the explainability of the model used is critical, especially in high-stakes areas such as medicine, finance, etc. Among the commonly used models, Linear Discriminant Analysis (LDA) is a widely used classification tool that is also explainable thanks to its ability to model class distributions and maximize class separation through linear feature combinations. Nevertheless, real-world data is frequently incomplete, presenting significant challenges for classification tasks and model explanations. In this paper, we propose a novel approach to LDA under missing data, termed Weighted missing Linear Discriminant Analysis (WLDA), to directly classify observations in data that contains missing values without imputation effectively by estimating the parameters directly on missing data and use a weight matrix for missing values to penalize missing entries during classification. Furthermore, we also analyze the theoretical properties and examine the explainability of the proposed technique in a comprehensive manner. Experimental results demonstrate that WLDA outperforms conventional methods by a significant margin, particularly in scenarios where missing values are present in both training and test sets.
@@
classification , missing data , parameter estimation , imputation
journal: Engineering Applications of Artificial Intelligence
[…]
[label1]organization=HCMC University of Technology and Education, city=Ho Chi Minh, country=Vietnam
[…]
[label4]organization=Simula Metropolitan Center for Digital Engineering, city=Oslo, country=Norway
@@ 1 Introduction
The above analysis motivates us to propose Weighted missing Linear Discriminant Analysis (WLDA) to address the challenges with LDA under missing data. More specifically, we want to address the dual challenges of handling missing data and ensuring model explainability in machine learning applications. The fundamental idea of the method is to use the DPER algorithm [6] for estimating the parameters, and the idea that when a value is missing, its contribution to classification should be zero. To do that, they utilize a weight matrix to penalize the contribution of a missing value. As will be illustrated experimentally, the proposed WLDA algorithm not only enhances the robustness and accuracy of models in the presence of missing data but also ensures that the decision-making process is transparent and comprehensible. This dual focus on data integrity and explainability is crucial for fostering trust, reliability, and broader acceptance of AI technologies in critical applications. Also, the proposed technique can be explained throughout from pre-modeling to post-modeling. In other words, we aim to adjust the LDA model to directly learn from missing data and give an explanation for this model. To achieve this, we use a weighted missing matrix to evaluate the contribution of each feature when it contains missing values. In summary, our contribution can be summarized as follows:
  1. 1.
    We introduce a novel algorithm, namely WLDA, capable of handling missing values in both training and test sets and treating each feature fairly without imputation,
  2. 2.
    We theoretically analyze WLDA properties,
  3. 3.
    We examine the explainability of the proposed algorithm in various aspects,
  4. 4.
    We conduct various experiments to demonstrate that our algorithm achieves superior performance compared to other methods under comparison.
  5. 5.
    We analyze the potential direction for future works.
@@ 3 Preliminary: Linear Discriminant Analysis
Linear Discriminant Analysis (LDA) assumes that the data from each class follows a multivariate Gaussian distribution and tries to minimize the total probability of misclassification [35]. Furthermore, the covariance matrices of all classes are assumed to be equal. The classification rule in LDA is to classify $\mathbf{x}$ to the $g^{th}$ class if
$\mathcal{L}_{g}(\mathbf{x})=\max\{\mathcal{L}_{1}(\mathbf{x}),\mathcal{L}_{2}(\mathbf{x}),...,\mathcal{L}_{G}(\mathbf{x})\}.$
Here,
@@
The decision boundary between two classes $g$ and $h$ is where their corresponding discriminant functions are equal, i.e.,
$\delta_{g}(\mathbf{x})=\delta_{h}(\mathbf{x}).$
Simplifying this equation, we get
$(\bm{\mu}^{(g)})^{T}\bm{\Sigma}^{-1}\mathbf{x}-\frac{1}{2}(\bm{\mu}^{(g)})^{T}\bm{\Sigma}^{-1}\bm{\mu}^{(g)}+\pi_{g}=(\bm{\mu}^{(h)})^{T}\bm{\Sigma}^{-1}\mathbf{x}-\frac{1}{2}(\bm{\mu}^{(h)})^{T}\bm{\Sigma}^{-1}\bm{\mu}^{(h)}+\pi_{h}.$
@@
$(\bm{\mu}^{(g)}-\bm{\mu}^{(h)})^{T}\bm{\Sigma}^{-1}\mathbf{x}+\frac{1}{2}\left((\bm{\mu}^{(h)})^{T}\bm{\Sigma}^{-1}\bm{\mu}^{(h)}-(\bm{\mu}^{(g)})^{T}\bm{\Sigma}^{-1}\bm{\mu}^{(g)}\right)+\log\frac{n_{g}}{n_{h}}=0.$
This equation represents a linear decision boundary. The weight vector $\mathbf{a}$ that defines the orientation of the decision boundary is given by
$\mathbf{a}_{gh}=\bm{\Sigma}^{-1}(\bm{\mu}^{(g)}-\bm{\mu}^{(h)}).$
The intercept term $a_{o}$ that defines the position of the decision boundary is given by
$a_{o}=\frac{1}{2}\left((\bm{\mu}^{(h)})^{T}\bm{\Sigma}^{-1}\bm{\mu}^{(h)}-(\bm{\mu}^{(g)})^{T}\bm{\Sigma}^{-1}\bm{\mu}^{(g)}\right)+\log\frac{n_{g}}{n_{h}}.$
Putting it all together, the decision boundary can be expressed in the linear form
$\mathbf{a}_{gh}^{T}\mathbf{x}+a_{0}=0.$
This linear equation defines the hyperplane that separates the two classes in the feature space.
@@ 4.3 Theoretical Analysis
$u_{o}=\frac{1}{2}\left((\bm{\mu}^{(h)})^{T}\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}}\bm{\mu}^{(h)}-(\bm{\mu}^{(g)})^{T}\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}}\bm{\mu}^{(g)}\right)+\log\frac{n_{g}}{n_{h}}.$
@@ 5.1 Correlation Visualization
  • 1.
    Correlation heatmap: This heatmap visually represents the correlation matrix between features. It uses a color gradient to show the strength and direction of correlations, typically ranging from -1 (strong negative correlation) to 1 (strong positive correlation), with 0 indicating no correlation. To convert a covariance matrix $\bm{\Sigma}$ to a correlation matrix $\bm{R}$, we use the following formula
    $\bm{R}=\bm{D}^{-1}\bm{\Sigma}\bm{D}^{-1}.$ (8)
    where $\bm{D}$ is a diagonal matrix containing the standard deviations of the features on the diagonal. The $i$-th diagonal element of $\bm{D}$ is $\sqrt{\sigma_{ii}}$, where $\sigma_{ii}$ is the variance of the $i$-th feature.
  • 2.
    Subtraction correlation heatmap: Also known as a difference heatmap, compares the ground truth correlation matrix $\bm{R}_{\text{true}}$ with an estimated correlation matrix $\bm{R}_{\text{est}}$. It visualizes where and to what extent the correlations between features differ between the two matrices
    $\Delta{\bm{R}}=\bm{R}_{\text{true}}-\bm{R}_{\text{est}}.$ (9)
    The elements $\Delta{r_{ij}}=r_{ij}^{\text{true}}-r_{ij}^{\text{est}}$ of $\Delta{\bm{R}}$ represents the directional differences in correlation coefficients $r_{ij}$ between the ground truth and estimated matrices. This heatmap emphasizes whether these differences are positive or negative, providing a clear indication of the direction and magnitude of variation. It aids in identifying areas where estimation techniques either overestimate or underestimate correlations compared to the ground truth.
  • 3.
    Local squared error correlation heatmap: The heatmap enhances the understanding of correlation estimation by visually depicting discrepancies between estimated correlation matrix $\bm{R}_{\text{est}}$ and true correlation matrix $\bm{R}_{\text{true}}$. The Squared Error is computed as
    $\Delta\bm{R}_{\text{squared}}=(\Delta\bm{R})\circ(\Delta\bm{R}).$ (10)
    By computing the element-wise squared differences, this heatmap provides a direct measure of the magnitude of error in correlation estimates across all pairs of features, aiding in method comparison and validation.
@@ 6.2.1 Classification performances
Datasets Missing rate Accuracy (mean $\pm$ standard deviation)
@@
Datasets Missing rate Accuracy (mean $\pm$ standard deviation)
@@ 6.2.2 Explainability analysis
(a) Correlation heatmap by using criteria (8)
(b) Local Squared Error correlation heatmap by using criteria (10)
(c) Subtraction correlation heatmap by using criteria (9)
@@ Appendix A Proof for the theorem 1
Theorem 1. The decision boundary between class $g$ and class $h$ of WLDA is given
$\mathbf{u}_{gh}^{T}\mathbf{x}+u_{o}=0,$
where
and
$u_{o}=\frac{1}{2}\left((\bm{\mu}^{(h)})^{T}\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}}\bm{\mu}^{(h)}-(\bm{\mu}^{(g)})^{T}\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}}\bm{\mu}^{(g)}\right)+\log\frac{n_{g}}{n_{h}}.$
Proof 1
@@
Therefore, the decision boundary between class $g$ and class $g$ is simplified to
$\mathbf{u}_{gh}^{T}\mathbf{x}+u_{o}=0.$
@@ Appendix B Proof for theorem 2
Theorem 2. Assume we have a data $\mathbf{X}$ is size of $n\times p$ which has $G$ classes and each class data $\mathbf{X}^{(g)}\sim\mathcal{N}(\bm{\mu}^{(g)},\bm{\Sigma})$ for $g=1,2,\dots,G$. For each instance $\mathbf{x}$, the expectation and variance of the WLDA classification function of $\mathbf{x}$ belongs to $g^{th}$ class given by
$\mathbb{E}[\mathcal{L}^{\mathbf{W}}_{g}(\mathbf{x})]=\pi_{g}-\frac{1}{2}\sum_{i=1}^{p}m_{i}^{2}w_{i}^{2},$
$\mathbf{Var}(\mathcal{L}^{\mathbf{W}}_{g}(\mathbf{x}))=\frac{1}{2}\sum_{i=1}^{p}m_{i}^{4}w_{i}^{4},$
$\mathbf{Bias}(\mathcal{L}^{\mathbf{W}}_{g}(\mathbf{x}))=\frac{1}{2}\left(p-\sum_{i=1}^{p}m_{i}^{2}w_{i}^{2}\right).$
@@
Note that $\bm{\mu}_{\mathbf{z}}=0$ and $\mathbf{A}=\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}}$ is a symmetric matrix, therefore we can apply the lemma 1 for $Q=\mathbf{z}^{T}(\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}})\mathbf{z}$
@@
Using the equation (13) in lemma 1 with the notice of $\bm{\mu}_{\mathbf{z}}=0$, we have
$\mathbf{Var}(Q)$ $=2\mathbf{Tr}((\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}})^{2}\bm{\Sigma}^{2})$
$=2\mathbf{Tr}(\mathbf{W}_{\mathbf{x}}^{4})$
$=2\sum_{i=1}^{p}m_{i}^{4}w_{i}^{4}.$
@@
This implies that for complete data
$\mathbb{E}[\mathcal{L}_{g}(\mathbf{x})]=\pi_{g}-\frac{1}{2}p.$
Therefore,
$\textbf{Bias}(\mathcal{L}^{\mathbf{W}}_{g}(\mathbf{x}))$ $=\left(\pi_{g}-\frac{1}{2}\sum_{i=1}^{p}m_{i}^{2}w_{i}^{2}\right)-\left(\pi_{g}-\frac{1}{2}p\right)$
$=\frac{1}{2}\left(p-\sum_{i=1}^{p}m_{i}^{2}w_{i}^{2}\right).$

### v3

# Directly Classification for missing Electronic Healthcare Records Data with Explainable Artificial Intelligence

@@ Abstract
As the adoption of Artificial Intelligence (AI) models expands into critical real-world applications, ensuring the explainability of these models becomes paramount, particularly in sensitive fields such as medicine and finance. Linear Discriminant Analysis (LDA) remains a popular choice for classification due to its interpretable nature, derived from its capacity to model class distributions and enhance class separation through linear combinations of features. However, real-world datasets often suffer from incomplete data, posing substantial challenges for both classification accuracy and model interpretability. In this paper, we introduce a novel and robust classification method, termed Weighted missing Linear Discriminant Analysis (WLDA), which extends LDA to handle datasets with missing values without the need for imputation. Our approach innovatively incorporates a weight matrix that penalizes missing entries, thereby refining parameter estimation directly on incomplete data. This methodology not only preserves the interpretability of LDA but also significantly enhances classification performance in scenarios plagued by missing data. We conduct an in-depth theoretical analysis to establish the properties of WLDA and thoroughly evaluate its explainability. Experimental results across various datasets demonstrate that WLDA consistently outperforms traditional methods, especially in challenging environments where missing values are prevalent in both training and test datasets. This advancement provides a critical tool for improving classification accuracy and maintaining model transparency in the face of incomplete data.
@@
classification; missing data; parameter estimation; imputation
@@ 1 Introduction
The above analysis motivates us to propose Weighted missing Linear Discriminant Analysis (WLDA) to address the challenges with LDA under missing data. More specifically, we want to address the dual challenges of handling missing data and ensuring model explainability in machine learning applications. The fundamental idea of the method is to use the DPER algorithm [6] for estimating the parameters, and the idea that when a value is missing, its contribution to classification should be zero. To do that, they utilize a weight matrix to penalize the contribution of a missing value. As will be illustrated experimentally, the proposed WLDA algorithm not only enhances the robustness and accuracy of models in the presence of missing data but also ensures that the decision-making process is transparent and comprehensible. This dual focus on data integrity and explainability is crucial for fostering trust, reliability, and broader acceptance of AI technologies in critical applications. Also, the proposed technique can be explained throughout from pre-modeling to post-modeling. In other words, we aim to adjust the LDA model to directly learn from missing data and give an explanation for this model. To achieve this, we use a weighted missing matrix to evaluate the contribution of each feature when it contains missing values. In summary, our contribution can be summarized as follows: (1) We introduce a novel algorithm, namely WLDA, capable of handling missing values in both training and test sets and treating each feature fairly without imputation; (ii) We theoretically analyze WLDA properties; (iii) We examine the explainability of the proposed algorithm in various aspects; (iv) We conduct various experiments to demonstrate that our algorithm achieves superior performance compared to other methods under comparison; (v) We analyze the potential direction for future works.
@@ 3 Preliminary: Linear Discriminant Analysis
Linear Discriminant Analysis (LDA) assumes that the data from each class follows a multivariate Gaussian distribution and tries to minimize the total probability of misclassification [35]. Furthermore, the covariance matrices of all classes are assumed to be equal. The classification rule in LDA is to classify $\mathbf{x}$ to the $g^{th}$ class if $\mathcal{L}_{g}(\mathbf{x})=\max\{\mathcal{L}_{1}(\mathbf{x}),\mathcal{L}_{2}(\mathbf{x}),...,\mathcal{L}_{G}(\mathbf{x})\}.$ Here,
@@
The decision boundary between two classes $g$ and $h$ is where their corresponding discriminant functions are equal, i.e., $\delta_{g}(\mathbf{x})=\delta_{h}(\mathbf{x}).$ Simplifying this equation, we get
@@
This equation represents a linear decision boundary. The weight vector $\mathbf{a}$ that defines the orientation of the decision boundary is given by $\mathbf{a}_{gh}=\bm{\Sigma}^{-1}(\bm{\mu}^{(g)}-\bm{\mu}^{(h)}).$
The intercept term $a_{o}$ that defines the position of the decision boundary is $a_{o}=\frac{1}{2}\left((\bm{\mu}^{(h)})^{T}\bm{\Sigma}^{-1}\bm{\mu}^{(h)}-(\bm{\mu}^{(g)})^{T}\bm{\Sigma}^{-1}\bm{\mu}^{(g)}\right)+\log\frac{n_{g}}{n_{h}}.$
Putting it all together, the decision boundary can be expressed in the linear form $\mathbf{a}_{gh}^{T}\mathbf{x}+a_{0}=0.$ This linear equation defines the hyperplane that separates the two classes in the feature space.
@@ 5.1 Correlation visualization
  • Correlation heatmap: This heatmap visually represents the correlation matrix between features. It uses a color gradient to show the strength and direction of correlations, typically ranging from -1 (strong negative correlation) to 1 (strong positive correlation), with 0 indicating no correlation. To convert a covariance matrix $\bm{\Sigma}$ to a correlation matrix $\bm{R}$, we use the following formula $\bm{R}=\bm{D}^{-1}\bm{\Sigma}\bm{D}^{-1}.$ where $\bm{D}$ is a diagonal matrix containing the standard deviations of the features on the diagonal.
  • Subtraction correlation heatmap: Also known as a difference heatmap, compares the ground truth correlation matrix $\bm{R}_{\text{true}}$ with an estimated correlation matrix $\bm{R}_{\text{est}}$. It visualizes where and to what extent the correlations between features differ between the two matrices $\Delta{\bm{R}}=\bm{R}_{\text{true}}-\bm{R}_{\text{est}}.$ The elements $\Delta{r_{ij}}=r_{ij}^{\text{true}}-r_{ij}^{\text{est}}$ of $\Delta{\bm{R}}$ represents the directional differences in correlation coefficients $r_{ij}$ between the ground truth and estimated matrices. This heatmap emphasizes whether these differences are positive or negative, providing a clear indication of the direction and magnitude of variation.
  • Local squared error correlation heatmap: The heatmap enhances the understanding of correlation estimation by visually depicting discrepancies between estimated correlation matrix $\bm{R}_{\text{est}}$ and true correlation matrix $\bm{R}_{\text{true}}$. The Squared Error is computed as $\Delta\bm{R}_{\text{squared}}=(\Delta\bm{R})\circ(\Delta\bm{R}).$
@@ 6.2.1 Classification performances
Dataset Missing rate Accuracy (mean $\pm$ standard deviation)
@@
Dataset Missing rate Accuracy (mean $\pm$ standard deviation)
@@ 6.2.2 Explainability analysis
(a) Correlation heatmap by using criteria (5.1)
(b) Local Squared Error correlation heatmap by using criteria (5.1)
(c) Subtraction correlation heatmap by using criteria (5.1)
@@ Appendix A Proof for the theorem 1
Theorem 1. The decision boundary between class $g$ and class $h$ of WLDA is given by $\mathbf{u}_{gh}^{T}\mathbf{x}+u_{o}=0,$ where $\mathbf{u}_{gh}=\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}}(\bm{\mu}^{(g)}-\bm{\mu}^{(h)}),$ and $u_{o}=\frac{1}{2}\left((\bm{\mu}^{(h)})^{T}\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}}\bm{\mu}^{(h)}-(\bm{\mu}^{(g)})^{T}\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}}\bm{\mu}^{(g)}\right)+\log\frac{n_{g}}{n_{h}}.$
Proof A.3.
@@
Therefore, the decision boundary between class $g$ and class $g$ is simplified to $\mathbf{u}_{gh}^{T}\mathbf{x}+u_{o}=0.$
@@ Appendix B Proof for theorem 2
Theorem 2. Assume we have a data $\mathbf{X}$ is size of $n\times p$ which has $G$ classes and each class data $\mathbf{X}^{(g)}\sim\mathcal{N}(\bm{\mu}^{(g)},\bm{\Sigma})$ for $g=1,2,\dots,G$. For each instance $\mathbf{x}$, the expectation and variance of the WLDA classification function of $\mathbf{x}$ belongs to $g^{th}$ class given by $\mathbb{E}[\mathcal{L}^{\mathbf{W}}_{g}(\mathbf{x})]=\pi_{g}-\frac{1}{2}\sum_{i=1}^{p}m_{i}^{2}w_{i}^{2},\mathbf{Var}(\mathcal{L}^{\mathbf{W}}_{g}(\mathbf{x}))=\frac{1}{2}\sum_{i=1}^{p}m_{i}^{4}w_{i}^{4},$ where $\mathbf{m}=(m_{1},m_{2},\dots,m_{p})$ and $\mathbf{w}=(w_{1},w_{2},\dots,w_{p})$ are the mask and weighted missing vector of $\mathbf{x}$ respectively. Moreover, the bias of this function is $\mathbf{Bias}(\mathcal{L}^{\mathbf{W}}_{g}(\mathbf{x}))=\frac{1}{2}\left(p-\sum_{i=1}^{p}m_{i}^{2}w_{i}^{2}\right).$
@@
Note that $\bm{\mu}_{\mathbf{z}}=0$ and $\mathbf{A}=\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}}$ is a symmetric matrix, therefore we can apply the lemma B.4 for $Q=\mathbf{z}^{T}(\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}})\mathbf{z}$
@@
Using the equation (10) in lemma B.4 with the notice of $\bm{\mu}_{\mathbf{z}}=0$, we have
$\mathbf{Var}(Q)$ $=2\mathbf{Tr}((\mathbf{W}_{\mathbf{x}}\bm{\Sigma}^{-1}\mathbf{W}_{\mathbf{x}})^{2}\bm{\Sigma}^{2})=2\mathbf{Tr}(\mathbf{W}_{\mathbf{x}}^{4})=2\sum_{i=1}^{p}m_{i}^{4}w_{i}^{4}.$
@@
This implies that for complete data $\mathbb{E}[\mathcal{L}_{g}(\mathbf{x})]=\pi_{g}-\frac{1}{2}p.$ Therefore, $\textbf{Bias}(\mathcal{L}^{\mathbf{W}}_{g}(\mathbf{x}))=\left(\pi_{g}-\frac{1}{2}\sum_{i=1}^{p}m_{i}^{2}w_{i}^{2}\right)-\left(\pi_{g}-\frac{1}{2}p\right)=\frac{1}{2}\left(p-\sum_{i=1}^{p}m_{i}^{2}w_{i}^{2}\right).$